# Photometric and spectroscopic variability of blue supergiant $\rho$ Leo

### v1

@@ 3.3 Inference of the inclination angle
Figure 13: Posterior distributions of the gravity-darkened physical parameters based on He I 4026 Å line profile modelling with zpektr using Hamiltonian Monte Carlo inference with STAN (HMC). The mean, the 16th and 84th percentile values for each parameter are shown in the histograms, where the blue curves indicate the joint probability density contours for the correlation between the pairs of parameters. The inclination angle that best match the data in average is ${21.7^{\circ}}^{+0.5}_{-0.4}$
Assuming the star to be a rigid rotator and that the stellar radiative flux in the outer atmospheric layers is anti-parallel to the effective gravity to preserve the flux/gravity ratio for all latitudes on the stellar surface, we used the zpektr code (Levenhagen2014ApJ...797...29L; refId0) to build a grid of $\approx$ 40 000 gravity-darkened model spectra using the Espinosa-Lara gravity-darkening prescription (Espinosa-Lara2011). The input plane-parallel model spectra to the code zpektr are hybrid LTE/NLTE set of model spectra that have been built with tlusty and synspec codes (Hubeny1988CoPhC..52..103H; Hubeny1995ApJ...439..875H) in NLTE regime (for models hotter than 15 000 K) and LTE Kurucz atmosphere models (Kurucz1979) with synspec (for models cooler than 15 000 K). The helium abundance was set to that of the star (0.2 dex), and the wavelength range was set to match that of the He I 4026 Å.
Figure 13 shows the results of the fittings of the observed He I 4026 Å  line profile with zpektr models using Hamiltonian Monte Carlo (HMC) inference with STAN (Stan-JSSv076i01; DUANE-HMC-1987216). For this, the HARPS spectrum observed on 12 February 2006 was used. The histograms and correlation plots, showing the joint probability density contours, reflect the correlations of each pair of parameters in the simulations. The best angle of inclination obtained in this analysis is ${21.7^{\circ}}^{+0.5}_{-0.4}$.
@@ 4 Discussion
In this study, we have obtained a wide range of periods from less than one day to up to $\sim$18 days repeating in all datasets (Tables in Section 3 and Table 7). The revealed pulsation pattern is more consistent with the theoretical predictions by 2013MNRAS.433.1246S for post-RSG blue supergiants, suggesting that $\rho$ Leo is on the blue loop of its evolution after the RSG stage.
Several periods identified from the TESS and K2 photometry, when considering the associated uncertainties, coincide with those found in the corresponding spectroscopic data. More importantly, a number of periods repeat in all four datasets, as illustrated in Table 7. Among them, the $\sim$17-day period consistently exhibits the strongest power, as shown in tables in Section 3. Also notable is the recurring $\sim$11-day period, even though its power is typically weaker. Both of these periods are independent and are not the result of combinations of other periods. We will explore these findings in more detail below.
Table 7: Summary table of detected periods. [^12] Table of repeating periods detected in spectroscopic and photometric datasets. For the spectral periods, the averaged values of three methods (GLS, GLSp and WWZ) are stated. Asterisk (*) indicates an uncertainty less than 0.04 days.
@@
First, we focus on the period of $\sim 17$ days. Figure 14 shows the K2 light curve and the first moment $\langle v_{\rm 1}\rangle$ of the 2017 spectroscopic data, both folded with a period of 17.212 days. As the observations were taken close in time, we analyse them together. This period gives a visually good fit to both curves, although they are slightly offset from each other. It is also only slightly shorter than the average period derived from the two datasets, 16.29 days.
An interesting feature is the shape of the spectroscopic phase curve – instead of a sinusoid, it has a sawtooth-like shape. This could have several possible interpretations. One is that it indicates radial pulsation. Alternatively, the sawtooth shape might suggest the presence of a close binary companion, with periodic mass exchange and possibly a circumbinary disk. These phenomena are consistent with theoretical models and observational findings reported by 1998AJ....116.1961H and 2024arXiv240416089P. Another possibility is that the sawtooth-like shape reflects mass loss modulated by some yet unknown process within the binary system (2008MNRAS.389.1605M).
Figure 14: Season 2017 first moment values (above) and K2 light curve (below) are phase folded with a period of 17.212 days. Sinusoidal fit to the photometric data is presented in red.
@@
In the literature, two other periods have been proposed as associated with stellar rotation. 2007AN....328.1170K report a period of 7.3 days. Our analysis of all datasets confirms the existence of this period. However, this period is not independent, but is a combination of other frequencies.
10.1093/mnras/sty308 reported a period of 26.8 days based on the combined K2 light curve and high-resolution HERMES spectroscopy, which was mostly obtained well before the K2 observations.
We recovered a similar period of approximately 25 days in the K2 data, consistent within uncertainties with their result, and our WWZ analysis confirms that this signal undergoes amplitude and phase modulation. However, we did not detect it in any of our spectroscopic datasets, and in the TESS photometry it appears only with very low amplitude, below the FAP threshold. This indicates that the reported 26.8-day signal is likely a transient feature rather than the true stellar rotation period.
A periodicity of about 5.5 days, detected across all our datasets, is close to the $f_{2}$ = 0.1956 d-1 ($P$ = 5.11 d) frequency reported by 10.1093/mnras/sty308 and is likely a harmonic of the 11-day mode. We do not confirm the presence of another frequency, $f_{3}$ = 0.4588 d-1 ($P$ = 2.18 d), which 10.1093/mnras/sty308 attributed to gravity waves: in the K2 data, the corresponding peak lies below the FAP threshold; a similar feature in the TESS photometry is likewise below FAP; and our GLSp analysis of the spectroscopic data from both 2017 and 2022 shows no significant power at this frequency.
@@ 5 Conclusions
    We identified set of repeating (quasi-)periods and their harmonics. The recurrence of most of these periods in different Seasons supports their authenticity. Several of these periods are detected using multiple methods (GLS, GLSp, and WWZ) and most appear in both photometric and spectroscopic data, further reinforcing their significance.
@@ Acknowledgements.
This work has used the ground-based research infrastructure of Tartu Observatory, funded through the projects TT8 (Estonian Research Council) and KosEST (EU Regional Development Fund). The authors acknowledge support from the Estonian Research Council grant IUT-40. This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No. 101079231 (EXOHOST), and from the United Kingdom Research and Innovation (UKRI) Horizon Europe Guarantee Scheme (grant number 10051045). This project has received funding from the European Union’s Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie Grant Agreement No. 823734 (POEMS) and is also Co-funded by the European Union (Project 101183150 – OCEANS). AA acknowledges support from the Estonian Research Council grant PRG 2159. Part of this work was supported by the German Deutsche Forschungsgemeinschaft, DFG project number Ts 17/2–1. Based on data obtained from the ESO Science Archive Facility with DOI https://doi.eso.org/10.18727/archive/33 under the ESO program 60.A-9036.
@@ Appendix A Photometric and spectroscopic time series
Figure 16: TESS photometry of $\rho$ Leo, observed from 6 November 2021 to 30 December 2021. Short- and long-term variability in brightness are clearly visible. The photometry was obtained just before the 2022 spectroscopic observations. Data available at CDS.
Figure 17: First moment (radial velocity) variations of the He I  6678.151 Å line profile during the Season 2017 (4 January 2017 – 16 May 2017). Each point corresponds to one separate spectrum. Typical error bar is shown in the top left corner. The 2017 spectra were obtained just before the K2 campaign. Data available at CDS.
Figure 18: Spectroscopic data in period 1 November 2021 to 15 May 2022 (so-called, Season 2022). Each point corresponds to one separate spectrum. In total – 1106 spectra, 39 nights. One preceded, two coincided with, and the rest closely followed the TESS photometry. Data available at CDS.
@@ Appendix B Log of spectroscopic observations
Table 8: Spectroscopic data for the Seasons 2017 and 2022.1313131477 spectra in total. SNR was measured near the He I line at 6678.151 Å. SNR and exposure time values are averaged per night.

### v2

@@ 3.3 Inference of the inclination angle
Figure 13: Posterior distributions of the gravity-darkened physical parameters based on He I 4026 Å line profile modelling with zpektr using Hamiltonian Monte Carlo inference with STAN (HMC). The mean and the 16th and 84th percentile values for each parameter are shown in the histograms, where the blue curves indicate the joint probability density contours for the correlation between the pairs of parameters. The inclination angle that best matches the data on average is ${21.7^{\circ}}^{+0.5}_{-0.4}$
Assuming the star to be a rigid rotator and that the stellar radiative flux in the outer atmospheric layers is anti-parallel to the effective gravity to preserve the flux/gravity ratio for all latitudes on the stellar surface, we used the zpektr code (Levenhagen2014ApJ...797...29L; refId0) to build a grid of $\approx$ 40 000 gravity-darkened model spectra using the Espinosa-Lara gravity-darkening prescription (Espinosa-Lara2011). The input plane-parallel model spectra to the code zpektr are a hybrid LTE/NLTE set of model spectra that were built with the tlusty and synspec codes (Hubeny1988CoPhC..52..103H; Hubeny1995ApJ...439..875H) in NLTE regime (for models hotter than 15 000 K) and LTE Kurucz atmosphere models (Kurucz1979) with synspec (for models cooler than 15 000 K). The helium abundance was set to that of the star (0.2 dex), and the wavelength range was set to match that of the He I 4026 Å.
Figure 13 shows the results of the fittings of the observed He I 4026 Å  line profile with the zpektr models using a Hamiltonian Monte Carlo (HMC) inference with STAN (Stan-JSSv076i01; DUANE-HMC-1987216). For this, the HARPS spectrum observed on 12 February 2006 was used. The histograms and correlation plots, showing the joint probability density contours, reflect the correlations of each pair of parameters in the simulations. The best angle of inclination obtained in this analysis is ${21.7^{\circ}}^{+0.5}_{-0.4}$.
@@ 4 Discussion
In this study, we have obtained a wide range of periods, from less than one day to $\sim$18 days repeating in all datasets (Tables in Section 3 and Table 7). The revealed pulsation pattern is more consistent with the theoretical predictions by 2013MNRAS.433.1246S for post-RSG blue supergiants, suggesting that $\rho$ Leo is on the blue loop of its evolution after the RSG stage.
Several periods identified from the TESS and K2 photometry, when considering the associated uncertainties, coincide with those found in the corresponding spectroscopic data. More importantly, a number of periods repeat in all four datasets, as illustrated in Table 7. Among them, the $\sim$17-day period consistently exhibits the strongest power, as shown in the tables in Section 3. Also notable is the recurring $\sim$11-day period, even though its power is typically weaker. Both of these periods are independent and are not the result of combinations of other periods. We explore these findings in more detail below.
Table 7: Summary table of detected periods. [^12] Table of repeating periods detected in the spectroscopic and photometric datasets. For the spectral periods, the averaged values of three methods (GLS, GLSp, and WWZ) are stated. An asterisk (*) indicates an uncertainty of less than 0.04 days.
@@
First, we focused on the period of $\sim 17$ days. Figure 14 shows the K2 light curve and the first moment $\langle v_{\rm 1}\rangle$ of the 2017 spectroscopic data, both folded with a period of 17.212 days. As the observations were taken close in time, we analysed them together. This period gives a visually good fit to both curves, although they are slightly offset from each other. It is also only slightly shorter than the average period derived from the two datasets, 16.29 days.
An interesting feature is the shape of the spectroscopic phase curve; instead of a sinusoid, it has a sawtooth-like shape. This could have several possible interpretations. One is that it indicates radial pulsation. Alternatively, the sawtooth shape might suggest the presence of a close binary companion, with periodic mass exchange and possibly a circumbinary disk. These phenomena are consistent with the theoretical models and observational findings reported by 1998AJ....116.1961H and 2024arXiv240416089P. Another possibility is that the sawtooth-like shape reflects mass loss modulated by some as-yet-unknown process within the binary system (2008MNRAS.389.1605M).
Figure 14: Season 2017 first-moment values (above) and K2 light curve (below) are phase-folded with a period of 17.212 days. The sinusoidal fit to the photometric data is presented in red.
@@
In the literature, two other periods are proposed as being associated with stellar rotation. 2007AN....328.1170K report a period of 7.3 days. Our analysis of all datasets confirms the existence of this period. However, this period is not independent, but is a combination of other frequencies. 10.1093/mnras/sty308 reported a period of 26.8 days based on the combined K2 light curve and high-resolution HERMES spectroscopy, which was mostly obtained well before the K2 observations.
We recovered a similar period of approximately 25 days in the K2 data, consistent within the uncertainties with their result, and our WWZ analysis confirms that this signal undergoes amplitude and phase modulation. However, we did not detect it in any of our spectroscopic datasets, and in the TESS photometry it appears only with very low amplitude, below the FAP threshold. This indicates that the reported 26.8-day signal is likely a transient feature rather than the true stellar rotation period.
A periodicity of about 5.5 days, detected across all our datasets, is close to the $f_{2}$ = 0.1956 d-1 ($P$ = 5.11 d) frequency reported by 10.1093/mnras/sty308 and is likely a harmonic of the 11-day mode. We do not confirm the presence of another frequency, $f_{3}$ = 0.4588 d-1 ($P$ = 2.18 d), which 10.1093/mnras/sty308 attributed to gravity waves. In the K2 data, the corresponding peak lies below the FAP threshold; a similar feature in the TESS photometry is likewise below FAP; and our GLSp analysis of the spectroscopic data from both 2017 and 2022 shows no significant power at this frequency.
@@ 5 Conclusions
    We identified a set of repeating (quasi-)periods and their harmonics. The recurrence of most of these periods in different seasons supports their authenticity. Several of these periods were detected using multiple methods (GLS, GLSp, and WWZ) and most appear in both photometric and spectroscopic data, further reinforcing their significance.
@@ Acknowledgements.
This work has used the ground-based research infrastructure of Tartu Observatory, funded through the projects TT8 (Estonian Research Council) and KosEST (EU Regional Development Fund). The authors acknowledge support from the Estonian Research Council grant IUT-40. This project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101079231 (EXOHOST), and from the United Kingdom Research and Innovation (UKRI) Horizon Europe Guarantee Scheme (grant number 10051045). This project has received funding from the European Union’s Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie Grant Agreement No. 823734 (POEMS) and is also Co-funded by the European Union (Project 101183150 – OCEANS). AA acknowledges support from the Estonian Research Council grant PRG 2159. Part of this work was supported by the German Deutsche Forschungsgemeinschaft, DFG project number Ts 17/2–1. Based on data obtained from the ESO Science Archive Facility with DOI https://doi.eso.org/10.18727/archive/33 under the ESO programme 60.A-9036.
@@ Appendix A Photometric and spectroscopic time series
Figure 16: TESS photometry of $\rho$ Leo, observed from 6 November 2021 to 30 December 2021. Short- and long-term variability in brightness are clearly visible. The photometry was obtained just before the 2022 spectroscopic observations. The data are available at CDS.
Figure 17: First-moment (radial velocity) variations of the He I  6678.151 Å line profile during Season 2017 (4 January 2017 – 16 May 2017). Each point corresponds to one separate spectrum. A typical error bar is shown in the top left corner. The 2017 spectra were obtained just before the K2 campaign. The data are available at CDS.
Figure 18: Spectroscopic data in the period 1 November 2021 to 15 May 2022 (Season 2022). Each point corresponds to one separate spectrum. In total there are 1106 spectra, 39 nights: one preceding, two coinciding with, and the rest closely following the TESS photometry. The data are available at CDS.
@@ Appendix B Log of spectroscopic observations
Table 8: Spectroscopic data for Seasons 2017 and 2022.1313131477 spectra in total. The S/N was measured near the He I line at 6678.151 Å. S/N and the exposure time values are averaged per night.